OpenStudy (anonymous):

Which ordered pair is a solution of the equation 4x − 2y = 16?

OpenStudy (anonymous):

A. (-5,-3) B. (-5,3) C. (6,-4) D. (6,4)

OpenStudy (anonymous):

(6,4): 4x - 2y = 4(6) - 2(4) = 24 - 8 = 16

OpenStudy (anonymous):

D